( n = 2.20) at an angle of 560 to the normal. What is the refracted angle?
1
Expert's answer
2019-02-17T09:37:07-0500

The Snell's law states

sin⁡〖θ_1 〗/sin⁡〖θ_2 〗 =n_2/n_1

So

sin⁡〖θ_2 〗=n_1/n_2 sin⁡〖θ_1 〗=1.00/2.20 〖 sin〗⁡〖56°〗=0.377

θ_2=sin^(-1)⁡0.377=22.1°